# On the contraction properties of Sinkhorn semigroups

### v1

@@ 2.2 Schrödinger bridges
starting from ${\cal P}_{0}$ with ${\cal S}_{0}={\cal Q}$, where, for a measure $\mu(dx)$ and a transition operator ${\cal S}(x,dy)$, the notation $(\mu\times{\cal S})^{\flat}$ denotes the ‘reverse’ measure
@@
${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}V_{2(n+1)}}$ $=$ ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}V_{2n+1}}:=V+\log{{\cal R}(e^{-U_{2n}})}$ (7)
@@
As shown in Section 6 of [2], under mild integrability assumptions the Schrödinger bridge distribution $\mathbb{P}$ solving (3) can also be computed from any Sinkhorn bridge distribution ${\cal P}_{2n}$, in the sense that for any $n\geq 0$ we have
@@
while Proposition 6.3 in [2] yields the formulae
@@
In addition, for any $n\geq 0$ we also have the commutation formulae
In the same vein, for any $n\geq 1$ we have
A proof for (11) and (12) is provided in Appendix E on page E.
@@ 2.3 Sinkhorn semigroups
$\pi_{2n}=\pi_{2p}\,{\cal S}^{\circ}_{2p,2n}\quad\mbox{\rm and}\quad{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\pi_{2n+1}}=\pi_{2p+1}\,{\cal S}^{\circ}_{2p+1,2n+1}.$
@@
with ${\cal S}^{\circ}_{2n,2n}=I={\cal S}^{\circ}_{2n+1,2n+1}$ the identity integral operator. The above framework shows that Sinkhorn semigroups belong to the class of time varying Markov processes (14) sharing the same invariant measures (13) at each individual step. Using the commutation properties (11) and (12) for any $n\geq p$ we readily check that
with the backward semigroups
By (11) and (12) we also obtain
@@ 3 Contraction of Markov operators
We recall some definitions of contraction coefficients for Markov operators, both in terms of the Hilbert projective metric and $\Phi$-entropies. These coefficients will be key to the analysis of contraction properties of Sinkhorn semigroups in Section 4.
@@ 3.1.1 Hilbert projective metric
[…]
@@ 3.1.2 Integral operator norms
${\left|\kern-1.07639pt\left|\kern-1.07639pt\left|\nu\right|\kern-1.07639pt\right|\kern-1.07639pt\right|}_{\varphi}$ ::=:= $\sup\{|\nu(f)|~{}:~{}\|f\|_{\varphi}\leq 1\}=\sup\{|\nu(f)|~{}:~{}\mbox{\rm osc}_{\varphi}(f)\leq 1\}$ (18)
$=$ $|\nu|(\varphi)\quad\mbox{\rm with the total variation}\quad|\nu|:=\nu_{+}+\nu_{-}.$
@@
For a detailed proof of the equivalent formulations in the latter definition (18), we refer to Proposition 8.2.16 in [38].
Whenever $\varphi\geq 1$, for any $\kappa>0$, we can construct the function $\phi_{\kappa}=1+\kappa\phi$ and obtain the norm equivalence formulae
$\kappa~{}\|\nu\|_{\sf tv}\leq\kappa~{}{\left|\kern-1.07639pt\left|\kern-1.07639pt\left|\nu\right|\kern-1.07639pt\right|\kern-1.07639pt\right|}_{\varphi}\leq{\left|\kern-1.07639pt\left|\kern-1.07639pt\left|\nu\right|\kern-1.07639pt\right|\kern-1.07639pt\right|}_{\varphi_{\kappa}}\leq(2+\kappa){\left|\kern-1.07639pt\left|\kern-1.07639pt\left|\nu\right|\kern-1.07639pt\right|\kern-1.07639pt\right|}_{\varphi}.$ (19)
The choice of condition $\varphi\geq 1/2$ in the definition of the $\varphi$-norm given in (18) is imposed only to recover the conventional total variation dual distance ${\left|\kern-1.07639pt\left|\kern-1.07639pt\left|\mu_{1}-\mu_{2}\right|\kern-1.07639pt\right|\kern-1.07639pt\right|}_{1/2}=\|\mu_{1}-\mu_{2}\|_{\sf tv}$ between probability measures $\mu_{1},\mu_{2}\in{\cal M}_{1}(\mathbb{X})={\cal M}_{1,\frac{1}{2}}(\mathbb{X})$, where ${\cal M}_{1,\varphi}(\mathbb{X})$ denotes the subset of probability measures $\nu$ s.t. $\nu(\varphi)<\infty$. Recall that for any $\mu_{1},\mu_{2}\in{\cal M}_{1}(S)$ and $\epsilon\in]0,1]$ we have
@@
3.1.3 Entropy and divergences
Consider a convex function
$\begin{array}[]{cccc}\Phi:&\mathbb{R}_{+}^{2}&\mapsto&\mathbb{R}\cup\{+\infty\}\\ &(u,v)&\leadsto&\Phi(u,v)\\ \end{array}$
such that for any $\alpha,v,w\in\mathbb{R}_{+}$ we have
$\Phi(\alpha w,\alpha v)=\alpha~{}\Phi(u,v)\quad\mbox{\rm and}\quad\Phi(1,1)=0.$ (21)
$D_{\Phi}(\nu_{1},\nu_{2}):=\int~{}\Phi\left(\frac{d\nu_{1}}{d\gamma}(x),\frac{d\nu_{2}}{d\gamma}(x)\right)~{}\gamma(dx).$
By homogeneity arguments, the above definition does not depend on the choice of the dominating measure $\gamma$. Note that
Whenever $\Phi(1,0)=\infty$, the $\Phi$-entropy coincides with the $\phi$-divergence in the sense of Csiszár [26], i.e., we also have
$\Phi(u,v)=\Phi_{1}(u,v):=u\log{(u/v)}\quad\mbox{\rm and}\quad\phi(u)=u\log{u}$
we recover the classical formulae
$\phi_{1}(u)=u\log{u}-(u-1).$
Other examples of $\phi$-divergences includes $\mathbb{L}_{p}$-norms, Havrda-Charvat entropies, $\alpha$-divergences, Jensen-Shannon divergence, squared Hellinger integrals and Kakutani-Hellinger, as well as Jeffreys and Rényi divergences [26, 37, 63, 65].

### v2

@@ 2.1 Sinkhorn bridges
starting from ${\cal P}_{0}$ with ${\cal S}_{0}={\cal Q}$, where, for a measure $\mu(dy)$ and a transition operator ${\cal S}(y,dx)$, the notation $(\mu\times{\cal S})^{\flat}$ denotes the ‘reverse’ measure
@@
$V_{2(n+1)}$ $=$ $V_{2n+1}:=V+\log{{\cal R}(e^{-U_{2n}})}.$ (18)
@@
Under mild integrability assumptions (see for instance Section 6 in [adm-24]) the Schrödinger bridge distribution $\mathbb{P}$ solving (4) can also be computed from any Sinkhorn bridge reference distribution ${\cal P}_{2n}$, in the sense that for any $n\geq 0$ we have
@@
while Proposition 6.3 in [adm-24] yields the formulae
@@ 2.2 Sinkhorn semigroups
$\pi_{2n}=\pi_{2p}\,{\cal S}^{\circ}_{2p,2n}\quad\mbox{\rm and}\quad\pi_{2n+1}=\pi_{2p+1}\,{\cal S}^{\circ}_{2p+1,2n+1}.$
@@
with ${\cal S}^{\circ}_{2n,2n}=I={\cal S}^{\circ}_{2n+1,2n+1}$ the identity integral operator. The above framework shows that Sinkhorn semigroups belong to the class of time varying Markov processes (23) sharing the same invariant measures (22) at each individual step.
Remark 2.1.
At every time step $n\geq 0$, the conjugate formulae (16) also show that ${\cal S}^{\circ}_{2n+1}:={\cal S}_{2n}{\cal S}_{2n+1}$ is the Markov transition of the simple alternating forward/backward Gibbs sampler with target measure
${\cal P}_{2n}(d(x,y))=\lambda_{U}(dx)~{\cal S}_{2n}(x,dy)=\pi_{2n}(dy)~{\cal S}_{2n+1}(y,dx)$
For $n=0$, the two-blocks Gibbs sampler with the forward/backward transition ${\cal S}^{\circ}_{1}:={\cal S}_{0}{\cal S}_{1}$ and target product measure ${\cal P}_{0}=\lambda_{U}\times{\cal S}_{0}$ can be interpreted as a sampling technique for $\lambda_{U}$ based on data augmentation or auxiliary variables (a.k.a. proximal sampling), see for instance [damlen1999gibbs, rendell2020global, vono2020asymptotically, vono2019split] and references therein. For a more detailed discussion on proximal samplers we also refer to the series of articles [mou, chewi-chen, chewi-phd, jiaming, jiaojiao, lee, guan]. Similarly, ${\cal S}^{\circ}_{2(n+1)}:={\cal S}_{2n+1}{\cal S}_{2(n+1)}$ is the Markov transition of the simple alternating forward/backward Gibbs sampler with target measure ${\cal P}_{2n+1}$.
2.3 Some stability theorems
In contrast with Kantorovich criteria or related Wasserstein-type semi-distances, even when $\mathbb{X}=\mathbb{Y}$ and the marginals coincide, $\lambda_{U}=\nu_{V}$, the identity-map deterministic coupling has infinite entropy. That is, for any reference measure ${\cal P}_{0}=(\lambda_{U}\times{\cal Q})$ of the form (3) we have
${\cal P}(d(x,y)):=\lambda_{U}(dx)~\delta_{x}(dy)\in\Pi_{U,V}\Longrightarrow\mbox{\rm Ent}({\cal P}~|~{\cal P}_{0})=+\infty.$
This fact, together with (5), shows that the Schrödinger bridge $\mathbb{P}$ and the convergence analysis of Sinkhorn bridges ${\cal P}_{2n}$ strongly depend on the form of the reference transition $\cal Q$} and the transition potential function $W$ discussed in (2).
Markov transitions of the form (2) with a bounded potential $W$ typically arise on finite or compact spaces. For instance, Markov processes on compact manifolds, such as diffusions on bounded manifolds reflected at the boundary, have Markov transitions of the form (2) with a uniformly bounded transition potential $W$ that depends on the time parameter [dpa, aronson, nash, varopoulos]. The convergence analysis of Sinkhorn bridges in this context is often developed using Hilbert Hilbert projective metrics [chen, deligiannidis, franklin, marino].
In the context of bounded transition potentials, the contraction principles of Markov operators presented in the article [dm-03] yield an alternative stability analysis that also applies to any generalized $\Phi$-entropies introduced in (13).
Theorem 2.2.
Assume that the transition potential $W$ is uniformly bounded. In this case, for any $\Phi$-entropy criteria $D_{\Phi}$ there is some parameters $\rho\in]0,1[$ and $c>0$ such that for any $n\geq 0$ we have the contraction inequalities
$D_{\Phi}\left(\pi_{2n},\nu_{V}\right)\leq\rho^{n}~D_{\Phi}\left(\pi_{0},\nu_{V}\right)\quad\mbox{and}\quad\left\|\log{({d\pi_{2n}}/{d\nu_{V}})}\right\|\leq c~\rho^{n}.$ (24)
The above inequalities remain valid when we replace $(\pi_{2n},\nu_{V})$ by $(\nu_{V},\pi_{2n})$. Similar estimates hold for $(\lambda_{U},\pi_{2n+1})$. There also exists $c>0$ such that for any $n\geq 0$ we have
$\mbox{\rm Ent}(\mathbb{P}~|~{\cal P}_{2n})\leq c~\rho^{n}.$ (25)
The estimates (24) are a consequence of the sandwich-type contraction inequalities stated in Theorem 4.1. A more refined version of the entropy estimate (25) in terms of Dobrushin contraction coeffcients is provided in Corollary 4.2
We underline that even on totally bounded state spaces such as $\mathbb{X}=\mathbb{Y}=]0,1[$ the transition potential function $W$ may be unbounded. Indeed, consider the reference Markov transition defined for any $x\in]0,1[$ by
${\cal Q}(x,dy)=\frac{1}{q(x)}~\left(1-|x-y|\right)~1_{]0,1[}(y)~dy\quad\mbox{\rm with}\quad\frac{1}{4}\leq q(x)=\frac{1}{2}-x(1-x)\leq\frac{1}{2}.$ (26)
In this context, we have
$W(x,y)=-\log{\left((x\wedge y)+(1-(x\vee y))\right)}+\log{q(x)}.$
The function $W(x,y)$ is locally bounded on $0,1[^{2}$ but tends to $\infty$ on boundaries. For instance, we obtain
$W\left(\frac{1}{n},1-\frac{1}{n}\right)\geq\log{\frac{n}{2}}-2\log{2}\stackrel{{\scriptstyle n\to\infty}}{{\longrightarrow}}+\infty.$
This rather simple model does not satisfy the boundedness condition of Theorem 2.2. To handle locally bounded transition potentials, consider the functions
${\cal U}_{\delta}:=\delta U-W^{V}\quad\mbox{\rm and}\quad{\cal V}_{\delta}:=\delta V-W_{U}$ (27)
with the integrated costs
Intuitively, the next theorem shows that Sinkhorn semigroups are stable as soon as the marginal-potentials $(U,V)$ dominate at infinity the (integrated) transition-potentials $(W_{U},W^{V})$. In terms of potential energy conversion, the marginal potentials should be large enough to compensate high energy transition losses. A more thorough discussion on these regularity conditions is provided in Section 4.4.
We formalize this physically intuitive principle with the second main Wasserstein-type stability theorem this article, which takes basically the following form.
Theorem 2.3.
Assume that $W$ is locally bounded and bounded below by a real number and $(U,V)$ are locally bounded with compact sub-level sets. In addition, there exists some $\delta\in]0,1[$ such that the functions $({\cal U}_{\delta/2},{\cal V}_{\delta/2})$ are locally bounded with compact sub-level sets and we have $\lambda_{(1-\delta)U}(1)\wedge\nu_{(1-\delta)V}(1)<\infty$. Then, there is some $\rho\in]0,1[$, $n_{0}\geq 1$ and $c>0$ s.t. for any $n\geq n_{0}$ we have
[…]
When $\psi$ satisfies the lower bound estimate (9) for some semi-metric $\varrho$ we also have the exponential Wasserstein estimates
$\|\pi_{2n}-\nu_{V}\|_{\tiny tv}\vee\mathbb{W}_{\varrho}(\pi_{2n},\nu_{V})\leq c~\rho^{n}.$ (30)
[…]
The estimates (29) are a consequence of Corollary 4.15 and the Wasserstein formulation (7) of weighted norms. The estimates (30) are a consequence of (10) and (29). More refined strict contraction estimates are provided in Corollary 4.14 and Corollary 4.16.
2.4 Illustrations and comments
Some illustrations
In Section 5, the stability theorems presented in Section 2.3 and Section 4 are illustrated in a variety of situations ranging from polynomial growth potentials (Section 5.1) and heavy tailed marginals on general normed spaces (Section 5.2), to more sophisticated boundary state space models (Section 5.3), semi-compact models including semi-discrete target spaces (Section 5.7) as well as finite mixture of statistical models (Section 5.9). Section 5.8 is dedicated to explicit descriptions and equivalent regularity conditions in the context of linear Gaussian models.
When $W$ is bounded, it does not appear useful to consider weighted total variation norms. In this context, the stability of Sinkhorn semigroups w.r.t. Hilbert’s metric is well known [borwein, chen, deligiannidis, franklin] and already yields exponential decays w.r.t. the total variation norm. On bounded subspaces $\mathbb{X}\subset\mathbb{R}^{d}$ the total variation norm also dominates Wasserstein distances w.r.t. the Euclidian norm but in general the target space $\mathbb{X}$ is non complete and then not Polish w.r.t. the Euclidean distance. In this context, it is desirable to find a complete distance on the target space that makes the corresponding Wasserstein distance in distribution space complete [bolley2008separability]. It is easy to construct examples of marginal models on non complete metric spaces. For instance, consider the marginal Beta distributions $\lambda_{U}=\nu_{V}$ on $\mathbb{X}=\mathbb{Y}=]0,1[$ equipped with the Euclidean metric $\varrho_{e}\left(x_{1},x_{2}\right)=|x_{1}-x_{2}|$ and defined by
$\nu_{V}(dx)=\frac{1}{6}~x(1-x)~\nu(dx)\quad\mbox{\rm with}\quad\nu(dx):=1_{]0,1[}(x)~dx.$
Applying (24) to the total variation distance (see (14)) we readily check the Wasserstein exponential decays
$\mathbb{W}_{\varrho_{e}}(\pi_{2n},\nu_{V})\leq\|\pi_{2n}-\nu_{V}\|_{\tiny tv}\leq\rho^{n}~\|\pi_{0}-\nu_{V}\|_{\tiny tv}.$
Note that the open unit interval $\mathbb{X}=]0,1[$ is not complete for the Euclidean metric but it is complete for the metric defined as
$\varrho_{c}\left(x_{1},x_{2}\right):=\left|\frac{1}{x_{1}}-\frac{1}{x_{2}}\right|+\left|\frac{1}{1-x_{1}}-\frac{1}{1-x_{2}}\right|\geq 2~{\varrho_{e}}(x_{1},x_{2}).$
On the other hand, for any parameter $\delta\in]0,1[$, if we have
$\psi(x):=\exp{(\delta V(x))}\geq\frac{6^{\delta}}{2}~\left(\frac{1}{x^{\delta}}+\frac{1}{(1-x)^{\delta}}\right)$
then
$\psi(x_{1})+\psi(x_{2})\geq\varrho_{c,\delta}\left(x_{1},x_{2}\right):=c_{\delta}~\varrho_{c}\left(x_{1},x_{2}\right)^{\delta}$
for some $c_{\delta}>0$. In this context, the l.h.s. estimate (29) yields the $\varrho_{c,\delta}$-Wasserstein estimates
[…]
as well as the weighted total variation norm decays
[…]
Theorem 2.3 also applies to Gaussian marginal distributions on $\mathbb{Y}=\mathbb{X}=\mathbb{R}^{d}$ and a transition potential $W$ with at most quadratic growth (cf. Section 5.1). In this context, applying (11) to $(p,q)=(2,1)$ we have the $2$-Wasserstein estimate (30), which implies
[…]
Section 5.8 provides an explicit description of the functions $({\cal U}_{\delta},{\cal V}_{\delta})$ in the context of linear-Gaussian models. For instance, for one-dimensional centered Gaussian marginals with unit variance and a symmetric Gaussian reference transition $\cal Q$} with variance $\tau>2$ we have
${2}/{\tau}<\delta<1\Longrightarrow{\cal U}_{\delta/2}(x)=c_{\delta}+\frac{x^{2}}{2}~\left(\frac{\delta}{2}-\frac{1}{\tau}\right)\stackrel{{\scriptstyle|x|\to\infty}}{{\longrightarrow}}+\infty\quad\mbox{\rm for some $c_{\delta}\in\mathbb{R}$. }$
A symmetric Gaussian reference transition $\cal Q$} with variance $\tau$ can be interpreted as the transition probability of a Brownian motion on a time horizon $\tau$. In this context, the exponential stability of Sinkhorn bridges w.r.t. the relative entropy for log-concave marginals at infinity, including Gaussian marginals, also holds for any time horizon $\tau>0$, see for instance [del2025stability, chiarini, durmus]. The article [adm-24] also provides Wasserstein exponential decays for any Gaussian marginals and any linear-Gaussian reference transition. For these classes of log-concave-type models, our Lyapunov approach yields new weighted total variation norm estimates for sufficiently large time horizons.
Energy barriers
Section 2.3 and Section 5 discuss a series of application models that can be handled using our Lyapunov approach. This section is concerned with some limitations of our approach and related important open research problems.
Firstly, note that Theorem 2.3 applies when $W$ is locally bounded. But even well defined absolutely continuous reference transitions on totally bounded state spaces such as $\mathbb{X}=\mathbb{Y}=]0,1[$ may have a potential $W$ which is not well defined on the whole space. Indeed, consider the reference Markov transition defined for any $x\in]0,1[$ by
$\begin{array}[]{l}\displaystyle{\cal Q}(x,dy)=\frac{1}{q(x)}~|x-y|~\nu(dy)\\ \\ \displaystyle\mbox{\rm with}\quad\nu(dy):=1_{]0,1[}(y)~dy\quad\mbox{\rm and}\quad\frac{1}{4}\leq q(x)=\frac{1}{2}-x(1-x)\leq\frac{1}{2}.\end{array}$
This rather elementary example is a a slight modification of (26) but in this case the function $W$ is not defined on the diagonal and for any $x\not=y$ we have
$\begin{array}[]{l}\displaystyle W(x,y)=-\log{|x-y|}+\log{q(x)}\Longrightarrow W(x,x)=+\infty.\end{array}$
To the best of our knowledge, the stability analysis of Sinkhorn bridges in the context of energy-barrier cost functions of this form remains an open research question. Absolutely continuous target marginals may also have an infinite potential, for instance
$\begin{array}[]{l}\displaystyle\lambda_{U}(dx):=4~|x-1/2|~\lambda(dx)\quad\mbox{\rm with}\quad\lambda(dx):=1_{]0,1[}(x)~dx\\ \\ \displaystyle\text{~~yields~~}U(x)=-\log{4}-\log{|x-1/2|}\Longrightarrow U(1/2)=+\infty.\end{array}$
In this case the potential function $U$ is not well defined on the whole state space, thus this marginal model cannot be written in terms of a Boltzmann-Gibbs measure of the form (1). Nevertheless for uniformly bounded transition potential functions, Theorem 2.2 applies to any class of marginal distributions.
Last but not least, note that the Weibull marginal distribution with shape parameter strictly less than one is locally bounded on $\mathbb{X}=]0,\infty[$ but does not have compact sub-level sets; for instance, we have
$\lambda_{U}(dx)=\frac{1}{2\sqrt{x}}~e^{-\sqrt{x}}~1_{]0,\infty[}(x)~dx\Longrightarrow U(x)=\sqrt{x}+\log{\sqrt{x}}+\log{2}\stackrel{{\scriptstyle x\to 0^{+}}}{{\longrightarrow}}-\infty.$
The potential function $U$ above is well defined on the whole state space but does not have compact sub-level sets. Even for bounded transition potential functions Theorem 2.3 cannot be used (but as mentioned above, Theorem 2.2 applies to any marginal models).
@@ 3 Contraction of Markov operators
Section 3.1 is dedicated to a brief description of the Hilbert (projective) metric and integral operator norms. Section 3.2 recalls some definitions of contraction coefficients for Markov operators in the context of Hilbert projective metrics, weighted total variation norms and $\Phi$-entropies. These coefficients will be key to the analysis of contraction properties of Sinkhorn semigroups in Section 4.
@@ 3.1.1 Hilbert projective metric
where $\|\cdot\|$ and $\|\cdot\|_{\sf tv}$ denote the supremum and total variation norms, respectively, and $H(\nu_{1},\nu_{2})=\infty$ when the measures $\nu_{1},\nu_{2}$ are not comparable. The proof of the r.h.s. estimate in (31) is provided in [atar] (cf. Lemma 1), see also [legland]. For a more thorough discussion on Hilbert projective metrics we refer the reader to the recent review and self contained article on Hilbert projective metrics [cohen].
@@ 3.1.2 Integral operator norms
[…]
$=$ $|\nu|(\varphi)\quad\mbox{\rm with the total variation measure}\quad|\nu|:=\nu_{+}+\nu_{-}.$
@@
For a detailed proof of the equivalent formulations in the latter definition (32), we refer to Proposition 8.2.16 in [penev]. The weighted Wasserstein formulation (7) of the $\varphi$-norm is a consequence of Kellerer’s extension of the Kantorovich-Rubinstein theorem [edwards2011kantorovich, kellerer1984duality, kellerer1982duality], see also the recent article [delgerber2025].
[…]
Recall that, given an arbitrary Polish space $S$, for any $\mu_{1},\mu_{2}\in{\cal M}_{1}(S)$ and $\epsilon\in]0,1]$ we have
@@
Choosing $\varphi\geq 1$ and $\kappa>0$, for any $\nu\in{\cal M}_{b}(\mathbb{X})$ we also have the norm equivalence formulae
[…]
For exponential functions of the form
$\varphi(x)=c~e^{g(x)}\mbox{\rm~~with some parameter $c\geq 0$ and~~}g(x_{1})+g(x_{2})\geq\varrho_{g}(x_{1},x_{2})$
for some semi-metric $\varrho_{g}$ we also have the estimate
$\varphi(x_{1})+\varphi(x_{2})\geq 2c~e^{g(x_{1})/2}e^{g(x_{2})/2}\geq\varrho_{\varphi}(x_{1},x_{2}):=2c~\left(e^{\frac{1}{2}\varrho_{g}(x_{1},x_{2})}-1\right).$
Also note that for any $p\geq 1$ we have
$\varrho_{\varphi}(x_{1},x_{2})\geq c_{p}~\varrho_{g}(x_{1},x_{2})^{p}\quad\mbox{\rm with}\quad c_{p}:=\frac{c}{2^{p-1}p!}$
and this implies that
[…]
Last but not least, on normed vector spaces consider a function of the form
$\varphi(x)\geq a+\|b(x)\|^{p}$
for some function $b(x)$ and $a,p>0$. In this case we have
$\varphi(x_{1})+\varphi(x_{2})\geq\varrho(x_{1},x_{2})\quad\mbox{\rm with}\quad\varrho(x_{1},x_{2}):=\frac{1}{2^{(p-1)_{+}}}~\|b(x_{1})-b(x_{2})\|^{p}$ (36)
as well as the Wasserstein estimate (10).